OpenStudy (anonymous):

Factor the following: 2x^2 - 128 @thewonderfuladele

OpenStudy (anonymous):

Im too tired to show the work lol, sorry 2(x+8)(x−8)

OpenStudy (solomonzelman):

1) factor out of 2. 2) use a^2-b^2=(a+b)(a-b) NOTE that in your case a=x and b=8.

OpenStudy (anonymous):

lol omg, you was suppose to help me lol

OpenStudy (anonymous):

I've been on here for an hour helping peeps so sorry :( I feel bad..

OpenStudy (anonymous):

lol, don't

OpenStudy (solomonzelman):

@cookiibabii93 I think I showed you what to do if you are one of those that want something else besides the answer.

OpenStudy (anonymous):

@SolomonZelman , thank you so much

OpenStudy (anonymous):

@SolomonZelman I know.. i was showing the work on the last couple questions he asked. im just getting tired of typing right now.. i guess im gnna have to log off soon.

OpenStudy (solomonzelman):

@thewonderfuladele if you were showing your work on similar questions and still needed to do this now, that means the user isn't taking anything out of the learning that you try to provide.

OpenStudy (anonymous):

I don't even need the work, just the answer, but I still want to learn it

OpenStudy (solomonzelman):

not the worse attitude...

OpenStudy (anonymous):

i could be wrong so dont want to answer but i think i have the answer

OpenStudy (anonymous):

correct me if im wrong √2xˆ2= √128 leading to 2x= 11.313 which is around x=5.66

OpenStudy (anonymous):

No, @KTA_Hype , I's suppose to look like the first answer. I'm factoring out the problems.

OpenStudy (anonymous):

i guess i just came at it from a different way

OpenStudy (anonymous):

earlier though u said u dont need the work just the answer so im confused